The projection on the line of axes are (6,2,3) then the lenght of line is
A. 11 
B. 7
C. 1
D. 0

Grade:12th pass

2 Answers

Saurabh Koranglekar
4 years ago
Dear student

Length of the line is sqrt( 36+4+9) = 7

Regards
Vikas TU
4 years ago
Dear student 
It can be written in vector form as 6i + 2j + 3k 
The distance of this line woud be equal to, 
sqrt(36+16+9) = sqrt(49) 
=> 7 cm